# Stability of ranks under field extensions

### v3

@@ 2.2. Properties of ranks
[20] For any $T\in\mathbb{K}^{n_{1}}\otimes\cdots\otimes\mathbb{K}^{n_{d}}$ and $S\in\mathbb{K}^{m_{1}}\otimes\cdots\otimes\mathbb{K}^{m_{d}}$, we have $\operatorname{SR}_{\mathbb{K}}(T\oplus S)=\operatorname{SR}_{\mathbb{K}}(T)+\operatorname{SR}_{\mathbb{K}}(S)$.
Lemma 2.2.
[32] For any $T\in\mathbb{F}_{q}^{n_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{n_{d}}$ and $1\leq k\leq d$, we have
$\operatorname{AR}_{\mathbb{F}_{q}}(T)=\sum_{i=1}^{d}n_{i}-n_{k}-\log_{q}\left(\left\lvert Z_{k}\left(T\right)\right\rvert\right).$
@@
[32] For any $T\in\mathbb{F}_{q}^{n_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{n_{d}}$ and $S\in\mathbb{F}_{q}^{m_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{m_{d}}$, we have $\operatorname{AR}_{\mathbb{F}_{q}}(T\oplus S)=\operatorname{AR}_{\mathbb{F}_{q}}(T)+\operatorname{AR}_{\mathbb{F}_{q}}(S)$.
3. Rank and subrank of a field extension
@@
$\varphi_{d}:\underbrace{\mathbb{F}\times\cdots\times\mathbb{F}}_{(d-1)\text{~{}copies}}\to\mathbb{F},\quad\varphi_{d}(x_{1},\dots,x_{d-1})=x_{1}\cdots x_{d-1}$ (7)
and denote the tensor corresponding to $\varphi_{d}$ by $T_{d,\mathbb{F}}\in\mathbb{F}^{\otimes d}$ where $\mathbb{F}^{\otimes d}\coloneqq\underbrace{\mathbb{F}\otimes\cdots\otimes\mathbb{F}}_{d\text{~{}copies}}$. The main purpose of this section is to estimate $R_{\mathbb{K}}(T_{d,\mathbb{F}})$ and $Q_{\mathbb{K}}(T_{d,\mathbb{F}})$ defined by (3).
@@
Suppose $\mathbb{K}$ is an infinite perfect field. For any integer $d\geq 2$ and any finite extension $\mathbb{F}/\mathbb{K}$, we have
$R_{\mathbb{K}}(T_{d,\mathbb{F}})\leq(d-1)(l-1)+1,\quad Q_{\mathbb{K}}(T_{d,\mathbb{F}})\geq\left\lceil\dfrac{l-1}{d-1}\right\rceil,$
where $l\coloneqq[\mathbb{F}:\mathbb{K}]$.
@@
Since $[\mathbb{F}:\mathbb{K}]=l<\infty$ and $\mathbb{K}$ is perfect, we have $\mathbb{F}=\mathbb{K}[\alpha]$ for some $\alpha\in\mathbb{F}$. Thus, we obtain the following commutative diagram:
@@
where $\Phi_{d}$ is the map multiplying $(d-1)$ polynomials of degree at most $(l-1)$ and $q$ is the restriction of the quotient map $\mathbb{K}[x]\to\mathbb{K}[\alpha]$. This implies that $R_{\mathbb{K}}(T_{d,\mathbb{F}})\leq R_{\mathbb{K}}(M_{d})$ where $M_{d}\in(\mathbb{K}^{l})^{\otimes(d-1)}\otimes\mathbb{K}^{(d-1)(l-1)+1}$ is the tensor corresponding to $\Phi_{d}$. Hence it is sufficient to estimate $R_{\mathbb{K}}(M_{d})$. To that end, we denote $N\coloneqq(d-1)(l-1)$ and pick distinct elements $a_{1},\dots,a_{N+1}$ in $\mathbb{K}$. For each $11\leq j\leq N+$ 1 and $1\leq k\leq d$, we define
$\eta_{jk}\coloneqq\begin{cases}\sum_{i=0}^{l-1}a_{j}^{i}\varepsilon_{i}&\text{~{}if~{}}1\leq k\leq d-1\\ \sum_{i=0}^{N}(V^{-1})_{ji}\varepsilon_{i}&\text{~{}if~{}}k=d\\ \end{cases}$
@@
and $\varepsilon_{s}:\mathbb{K}[x]\to\mathbb{K}$ is the linear function such that $\varepsilon_{s}(x^{t})=\delta_{st}$, $0\leq s,t$. We consider
@@
Given $f_{1},\dots,f_{d-1}\in\mathbb{K}[x]_{\leq l-1}$, it is clear that $M^{\prime}_{d}(f_{1},\dots,f_{d-1})$ computes the product polynomial $\prod_{k=1}^{d}f_{k}\in\mathbb{K}[x]_{\leq N}$ by interpolation at points $a_{1},\dots,a_{N+1}\in\mathbb{K}$. Thus $R_{\mathbb{K}}(M_{d})=R_{\mathbb{K}}(M^{\prime}_{d})\leq N+1$.
For $Q_{\mathbb{K}}(T_{d,\mathbb{F}})$, we let $m\coloneqq\lfloor(l-1)/(d-1)\rfloor$ and consider the commutative diagram:
@@
where $\Psi_{d}$ denotes the map multiplying $(d-1)$ polynomials of degree at most $m$. This implies that $Q_{\mathbb{K}}(T_{d,\mathbb{F}})\geq Q_{\mathbb{K}}(\Psi_{d})$. For distinct $b_{1},\cdots,b_{m+1}\in\mathbb{K}$, we define a linear map
@@
[…]
where $\psi_{d}$ is defined as $\psi_{d}(x_{1},\dots,x_{d-1})\coloneqq x_{1}\cdots x_{d-1}$ for the trivial $\mathbb{K}$-algebra $\mathbb{K}^{m+1}$. Since the tensor corresponding to $\psi_{d}$ is the identity tensor $\operatorname{Id}_{m+1}$, we obtain $Q_{\mathbb{K}}(T_{d,\mathbb{F}})\geq Q_{\mathbb{K}}(\Psi_{d})\geq m+1$. ∎
If $\mathbb{K}=\mathbb{F}_{q}$ is a finite field, we record the following estimate of $\operatorname{R}_{\mathbb{F}_{q}}(T_{d,\mathbb{F}})$ for easy reference.
@@
[3, Theorem 3] Given any integer $d>0$ and prime power $q$, there is a constant $C\coloneqq C(d,q)>0$ such that for any finite field extension $\mathbb{F}/\mathbb{F}_{q}$, we have $\operatorname{R}_{\mathbb{F}_{q}}(T_{d,\mathbb{F}})\leq Cl$ where $l\coloneqq[\mathbb{F}:\mathbb{F}_{q}]$.
We remark that if $d=3$, Theorem 3.2 reduces to a classical result in algebraic complexity theory [30, Theorem 2]. Next we consider $\operatorname{Q}_{\mathbb{K}}(T_{d,\mathbb{F}})$ when $\mathbb{K}$ is a finite field. We borrow the idea behind the Chudnovsky-Chudnovsky method [7, 39, 2, 3] designed to obtain a sharp upper bound of $\operatorname{R}_{\mathbb{K}}(T_{3,\mathbb{F}})$. To achieve the goal, we need to establish some technical lemmas.
@@
Let $\mathbb{F}/\mathbb{F}_{q}$ be an algebraic function field with genus $g$ and let $N,d,n$ be positive integers such that $N\geq g+1$, $(d-1)(N+g-1)<n$. If $\mathbb{F}/\mathbb{F}_{q}$ has at least $N$ degree one places $\mathfrak{m}_{1},\dots,\mathfrak{m}_{N}$ and a degree $n$ place $\mathfrak{m}$, then there exists a divisor $D$ such that
@@
Let $n,d,N$ be positive integers. Suppose $\mathbb{F}/\mathbb{F}_{q}$ is an algebraic function field of genus $g$ such that
@@
Then $\operatorname{Q}_{\mathbb{F}_{q}}(T_{d,\mathbb{F}_{q^{n}}})\geq N$.
@@
[…]
Here $\widetilde{S}:\mathcal{O}\to\mathcal{O}/{\mathfrak{m}}\simeq\mathbb{F}_{q^{N}}$ and $\widetilde{E}:\mathcal{O}_{1}\bigcap\cdots\bigcap\mathcal{O}_{N}\to\mathbb{F}_{q}^{N}$ are evaluation maps defined by $\widetilde{S}(f)=f(\mathfrak{m})$ and $\widetilde{E}(f)=(f(\mathfrak{m}_{1}),\dots,f(\mathfrak{m}_{N}))$, respectively. We claim that vertical maps in (12) are well-defined and the diagram commutes. This implies that
$\operatorname{Q}_{\mathbb{F}_{q}}(T_{d,\mathbb{F}_{q^{n}}})=\operatorname{Q}_{\mathbb{F}_{q}}(T_{d,\mathcal{O}/\mathfrak{m}})\geq\operatorname{Q}_{\mathbb{F}_{q}}(T_{d,\mathbb{F}_{q}^{N}})=N.$
It is left to prove the claim. We first verify that vertical maps are well-defined. According to Lemma 3.3, we have $v_{\mathfrak{m}}(D)=v_{\mathfrak{m}_{1}}(D)=\cdots=v_{\mathfrak{m}_{N}}(D)$. This implies that $v_{\mathfrak{m}}(f)\geq 0$ if $f\in\mathcal{L}(D)$ and $v_{\mathfrak{m}_{j}}(f)\geq 0$ for all $1\leq j\leq N$ if $f\in\mathcal{L}((d-1)D)$, which lead to $\mathcal{L}(D)\subseteq\mathcal{O}$ and $\mathcal{L}((d-1)D)\subseteq\bigcap_{j=1}^{N}\mathcal{O}_{j}$. Therefore, maps $\widetilde{S}\circ E^{-1}$ and $\widetilde{E}\circ S^{-1}$ are well-defined.
Next, we prove the commutativity of (12). For any $a_{1},\dots,a_{d-1}\in\mathbb{F}_{q}^{N}$, we denote $f_{j}\coloneqq E^{-1}(a_{j})\in\mathscr{L}(D)$, $11\leq j\leq d-$ 1. Thus
@@
$\widetilde{E}\left(\prod_{j=1}^{d-1}f_{j}\right)=\prod_{j=1}^{d-1}\widetilde{E}(f_{j})=\prod_{j=1}^{d-1}a_{i}=\varphi_{d}(a_{1},\dots,a_{d-1}).\qed$
@@
If $(l-1)\geq(d-1)(n-1)$, then we have $\operatorname{Q}_{\mathbb{F}_{q}}(T_{d,\mathbb{F}_{q^{n}}})\leq\operatorname{Q}_{\mathbb{F}_{q}}(T_{d,\mathbb{F}_{q^{l}}})$.
@@
Since $\mathbb{F}_{q}$ is a perfect field, we observe that (8) (resp. (9)) holds for $(\mathbb{K},\mathbb{F})=(\mathbb{F}_{q},\mathbb{F}_{q^{n}})$ (resp. $(\mathbb{K},\mathbb{F})=(\mathbb{F}_{q},\mathbb{F}_{q^{l}})$). Thus, we have $T_{d,\mathbb{F}_{q^{n}}}\preceq M_{d,n,\mathbb{F}_{q}}\preceq T_{d,\mathbb{F}_{q^{l}}}$, from which we obtain $Q_{\mathbb{F}_{q}}(T_{d,\mathbb{F}_{q^{n}}})\leq Q_{\mathbb{F}_{q}}(T_{d,\mathbb{F}_{q^{l}}})$. Here $M_{d,n,\mathbb{F}_{q}}$ is the tensor corresponding to the multilinear map multiplying $(d-1)$ $\mathbb{F}_{q}$-polynomials of degree at most $(n-1)$:
$\Phi_{d}:\underbrace{\mathbb{F}_{q}[x]_{\leq n-1}\times\cdots\times\mathbb{F}_{q}[x]_{\leq n-1}}_{(d-1)\text{~{}copies}}\to\mathbb{F}_{q}[x]_{(d-1)(n-1)},\quad\Phi_{d}(f_{1},\dots,f_{d-1})=\prod_{j=1}^{d-1}f_{j}.\qed$
@@
$\operatorname{Q}_{\mathbb{F}_{q}}(T_{d,\mathbb{F}_{q^{mn}}})\geq\operatorname{Q}_{\mathbb{F}_{q}}(T_{d,\mathbb{F}_{q^{m}}})\operatorname{Q}_{\mathbb{F}_{q^{m}}}(T_{d,\mathbb{F}_{q^{mn}}})$.
@@
$T_{d,\mathbb{F}_{q^{m}}}\in\underbrace{\mathbb{F}_{q^{m}}\otimes_{\mathbb{F}_{q}}\cdots\otimes_{\mathbb{F}_{q}}\mathbb{F}_{q^{m}}}_{d\text{~{}copies}},\quad T_{d,\mathbb{F}_{q^{mn}}}\in\underbrace{\mathbb{F}_{q^{mn}}\otimes_{\mathbb{F}_{q^{m}}}\cdots\otimes_{\mathbb{F}_{q^{m}}}\mathbb{F}_{q^{mn}}}_{d\text{~{}copies}}$
@@
$T_{d,\mathbb{F}_{q^{m}}}\boxtimes_{\mathbb{F}^{m}_{q}}T_{d,\mathbb{F}_{q^{mn}}}=T_{d,\mathbb{F}_{q^{mn}}}\in\underbrace{\mathbb{F}_{q^{mn}}\otimes_{\mathbb{F}_{q}}\cdots\otimes_{\mathbb{F}_{q}}\mathbb{F}_{q^{mn}}}_{d\text{~{}copies}},$
The desired inequality follows immediately from the definition of the subrank. ∎
@@
For any positive integers $d,n$ and prime power $q$, we have $\operatorname{Q}_{\mathbb{F}_{q}}(T_{d,\mathbb{F}_{q^{n}}})\geq\frac{n}{48d^{2}q}$.
@@
We denote $\sigma_{q}(n)\coloneqq Q_{\mathbb{F}_{q}}(T_{d,\mathbb{F}_{q^{n}}})$ and we first estimate $\sigma_{q^{2}}(n)$. For each positive integer $k$, we let $g_{k}$ and $N_{k}$ (if $k\geq 3$) be positive integers defined as in Lemma A.6. We consider
@@
By (30), we have $N_{m}\geq g_{m}$. The maximality of $m$ and $N$ leads to

### v4

@@ 2.2 Properties of ranks
[24] For any $T\in\mathbb{K}^{n_{1}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{n_{d}}$ and $S\in\mathbb{K}^{m_{1}}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{K}^{m_{d}}$, we have $\operatorname{SR}_{\mathbb{K}}(T\oplus S)=\operatorname{SR}_{\mathbb{K}}(T)+\operatorname{SR}_{\mathbb{K}}(S)$.
Lemma $2.$
[38] For any $T\in\mathbb{F}_{q}^{n_{1}}\otimes_{\mathbb{F}_{q}}\cdots\otimes_{\mathbb{F}_{q}}\mathbb{F}_{q}^{n_{d}}$ and $1\leq k\leq d$, we have
[…]
@@
[38] For any $T\in\mathbb{F}_{q}^{n_{1}}\otimes_{\mathbb{F}_{q}}\cdots\otimes_{\mathbb{F}_{q}}\mathbb{F}_{q}^{n_{d}}$ and $S\in\mathbb{F}_{q}^{m_{1}}\otimes_{\mathbb{F}_{q}}\cdots\otimes_{\mathbb{F}_{q}}\mathbb{F}_{q}^{m_{d}}$, we have $\operatorname{AR}_{\mathbb{F}_{q}}(T\oplus S)=\operatorname{AR}_{\mathbb{F}_{q}}(T)+\operatorname{AR}_{\mathbb{F}_{q}}(S)$.
The following lemma is a well-known fact concerning the monotonicity of analytic rank with respect the restriction.
Lemma 2.4 (Monotonicity of analytic rank).
[38, 9] Given $T\in\mathbb{F}_{q}^{n_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{n_{d}}$ and $S\in\mathbb{F}_{q}^{m_{1}}\otimes\cdots\otimes\mathbb{F}_{q}^{m_{d}}$ such that $T\preceq S$, we have $\operatorname{AR}_{\mathbb{F}_{q}}(T)\leq\operatorname{AR}_{\mathbb{F}_{q}}(S)$.
3 Rank and subrank of a field extension
@@
$\varphi_{d}:\underbrace{\mathbb{F}\times\cdots\times\mathbb{F}}_{(d-1)\text{~copies}}\to\mathbb{F},\quad\varphi_{d}(x_{1},\dots,x_{d-1})=x_{1}\cdots x_{d-1}.$ (7)
As a $\mathbb{K}$-multilinear map, $\varphi_{d}$ naturally corresponds to a tensor $\widetilde{M}_{d,\mathbb{F}}\in\underbrace{\mathbb{F}^{\ast}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{F}^{\ast}}_{(d-1)\text{~copies}}\otimes_{\mathbb{K}}\mathbb{F}$. By choosing and fixing a $\mathbb{K}$-basis of $\mathbb{F}$, we may identify $\mathbb{F}^{\ast}$ with $\mathbb{F}$ to obtain a tensor $M_{d,\mathbb{F}}\in\underbrace{\mathbb{F}\otimes_{\mathbb{K}}\cdots\otimes_{\mathbb{K}}\mathbb{F}}_{d\text{~copies}}$. Although $M_{d,\mathbb{F}}$ depends on the choice of $\mathbb{K}$-basis of $\mathbb{F}$, the ranks of $M_{d,\mathbb{F}}$ discussed in this paper do not. In fact, the ranks of $M_{d,\mathbb{F}}$, $\widetilde{T}_{d,\mathbb{F}}$ and $\varphi_{d}$ are all equal. We identify $\mathbb{F}^{\ast}$ with $\mathbb{F}$ just to simplify notations. All our arguments can be adjusted accordingly for $\widetilde{T}_{d,\mathbb{F}}$ and $\varphi_{d}$. The main purpose of this section is to estimate $\operatorname{R}_{\mathbb{K}}(M_{d,\mathbb{F}})$ and $\operatorname{Q}_{\mathbb{K}}(M_{d,\mathbb{F}})$ defined by (3).
@@
Let $\mathbb{F}$ be a finite field extension of $\mathbb{K}$ and let $l\coloneqq[\mathbb{F}:\mathbb{K}]$. Suppose $d\geq 2$ is an integer.
  1. (a)
    If $\mathbb{K}$ is either an infinite perfect field or a finite field with $\lvert\mathbb{K}\rvert\geq(d-1)(l-1)+1$, then
    $\operatorname{R}_{\mathbb{K}}(M_{d,\mathbb{F}})\leq(d-1)(l-1)+1.$
  2. (b)
    If $\mathbb{K}$ is either an infinite perfect field or a finite field, then
    $\operatorname{Q}_{\mathbb{K}}(M_{d,\mathbb{F}})\geq\max\left\{\left\lceil\dfrac{l-1}{d-1}\right\rceil,\lvert\mathbb{K}\rvert\right\}$.
@@
Assume that $\mathbb{K}$ is infinite and perfect. We have $\mathbb{F}=\mathbb{K}[\alpha]$ for some $\alpha\in\mathbb{F}$. Thus, we obtain the following commutative diagram:
@@
where $\Phi_{d}$ is the map multiplying (d1)(d-1) polynomials of degree at most (l1)(l-1) and $q$ is the restriction of the quotient map $\mathbb{K}[x]\to\mathbb{K}[\alpha]$. This implies that $\operatorname{R}_{\mathbb{K}}(M_{d,\mathbb{F}})\leq\operatorname{R}_{\mathbb{K}}(M_{d})$ where $M_{d}\in(\mathbb{K}^{l})^{\otimes(d-1)}\otimes\mathbb{K}^{(d-1)(l-1)+1}$ is the tensor corresponding to $\Phi_{d}$. Hence it is sufficient to estimate $\operatorname{R}_{\mathbb{K}}(M_{d})$. To that end, we denote $N\coloneqq(d-1)(l-1)$ and pick distinct elements $a_{1},\dots,a_{N+1}$ in $\mathbb{K}$. For each $1\leq j\leq N+1$ and $1\leq k\leq d$, we define
[…]
@@
and $\varepsilon_{s}:\mathbb{K}[x]\to\mathbb{K}$ is the linear function such that $\varepsilon_{s}(x^{t})=\delta_{st}$, $0\leq s,t\leq N$. We consider
@@
Given $f_{1},\dots,f_{d-1}\in\mathbb{K}[x]_{\leq l-1}$, it is clear that $M^{\prime}_{d}(f_{1},\dots,f_{d-1})$ computes the product polynomial $\prod_{k=1}^{d-1}f_{k}\in\mathbb{K}[x]_{\leq N}$ by interpolation at points $a_{1},\dots,a_{N+1}\in\mathbb{K}$. Thus $\operatorname{R}_{\mathbb{K}}(M_{d})=\operatorname{R}_{\mathbb{K}}(M^{\prime}_{d})\leq N+1$.
For $\operatorname{Q}_{\mathbb{K}}(M_{d,\mathbb{F}})$, we let $m\coloneqq\lfloor(l-1)/(d-1)\rfloor$ and consider the commutative diagram:
@@
where $\Psi_{d}$ denotes the map multiplying (d1)(d-1) polynomials of degree at most $m$. This implies that $\operatorname{Q}_{\mathbb{K}}(M_{d,\mathbb{F}})\geq\operatorname{Q}_{\mathbb{K}}(\Psi_{d})$. For distinct $b_{1},\cdots,b_{m+1}\in\mathbb{K}$, we define a linear map
@@
$\underbrace{\mathbb{K}[x]_{\leq m}\times\cdots\times\mathbb{K}[x]_{\leq m}}_{(d-1)\text{~copies}}$${\mathbb{K}[x]}_{\leq l-1}$$\underbrace{\mathbb{K}^{m+1}\times\cdots\times\mathbb{K}^{m+1}}_{(d-1)\text{~copies}}$$\mathbb{K}^{m+1}$$\scriptstyle{\Psi_{d}}$$\scriptstyle{S}$$\scriptstyle{S^{-1}\times\cdots\times S^{-1}}$$\scriptstyle{\psi_{d}}$
where $\psi_{d}$ is defined as $\psi_{d}(x_{1},\dots,x_{d-1})\coloneqq x_{1}\cdots x_{d-1}$ for the trivial $\mathbb{K}$-algebra $\mathbb{K}^{m+1}$. Since the tensor corresponding to $\psi_{d}$ is the identity tensor $\operatorname{Id}_{m+1}$, we obtain $\operatorname{Q}_{\mathbb{K}}(M_{d,\mathbb{F}})\geq\operatorname{Q}_{\mathbb{K}}(\Psi_{d})\geq m+1$.
When $\mathbb{K}$ is a finite field with $|\mathbb{K}|\geq(d-1)(l-1)+1$ (resp. $|\mathbb{K}|\geq\lfloor(l-1)/(d-1)\rfloor+1$), the above argument for $\operatorname{R}_{\mathbb{K}}(M_{d,\mathbb{F}})$ (resp. $\operatorname{Q}_{\mathbb{K}}(M_{d,\mathbb{F}})$) applies verbatim. For $|\mathbb{K}|\leq\lfloor(l-1)/(d-1)\rfloor$, we set $m\coloneqq|\mathbb{K}|$ in the argument for $\operatorname{Q}_{\mathbb{K}}(M_{d,\mathbb{F}})$ to ensure the interpolation remains applicable. ∎
If $\mathbb{K}=\mathbb{F}_{q}$ is a finite field, we record the following estimate of $\operatorname{R}_{\mathbb{F}_{q}}(M_{d,\mathbb{F}})$ for easy reference.
@@
[5, Theorem 3] Given any integer $d>0$ and prime power $q$, there is a constant $C\coloneqq C(d,q)>0$ such that for any finite field extension $\mathbb{F}/\mathbb{F}_{q}$, we have $\operatorname{R}_{\mathbb{F}_{q}}(M_{d,\mathbb{F}})\leq Cl$ where $l\coloneqq[\mathbb{F}:\mathbb{F}_{q}]$.
We remark that if $d=3$, Theorem 3.2 improves a classical result in algebraic complexity theory [37, Theorem 2]. Next we consider $\operatorname{Q}_{\mathbb{K}}(M_{d,\mathbb{F}})$ when $\mathbb{K}$ is a finite field. We borrow the idea behind the Chudnovsky-Chudnovsky method [11, 45, 4, 5] designed to obtain a sharp upper bound of $\operatorname{R}_{\mathbb{K}}(M_{3,\mathbb{F}})$. To achieve the goal, we need to establish some technical lemmas.
@@
Let $\mathbb{F}/\mathbb{F}_{q}$ be an algebraic function field with genus $g$ and let $N,d,n$ be positive integers such that $d\geq 2$, $N\geq g+1$, (d1)(N+g1)<n(d-1)(N+g-1)<n. If $\mathbb{F}/\mathbb{F}_{q}$ has at least $N$ degree one places $\mathfrak{m}_{1},\dots,\mathfrak{m}_{N}$ and a degree $n$ place $\mathfrak{m}$, then there exists a divisor $D$ such that
@@
Let $n,d\geq 2,N$ be positive integers. Suppose $\mathbb{F}/\mathbb{F}_{q}$ is an algebraic function field of genus $g$ such that
@@
Then $\operatorname{Q}_{\mathbb{F}_{q}}(M_{d,\mathbb{F}_{q^{n}}})\geq N$.
@@
$\underbrace{\mathbb{F}_{q}^{N}\times\cdots\times\mathbb{F}_{q}^{N}}_{(d-1)\text{~copies}}$$\mathbb{F}_{q}^{N}$$\underbrace{\mathcal{O}/\mathfrak{m}\times\cdots\times\mathcal{O}/\mathfrak{m}}_{(d-1)\text{~copies}}$$\mathcal{O}/\mathfrak{m}$$\scriptstyle{\varphi_{d}}$$\scriptstyle{(\widetilde{S}\circ E^{-1})\times\cdots\times(\widetilde{S}\circ E^{-1})}$$\scriptstyle{\psi_{d}}$$\scriptstyle{\widetilde{E}\circ S^{-1}}$ (12)
Here $\widetilde{S}:\mathcal{O}\to\mathcal{O}/{\mathfrak{m}}\simeq\mathbb{F}_{q^{N}}$ and $\widetilde{E}:\mathcal{O}_{1}\cap\cdots\cap\mathcal{O}_{N}\to\mathbb{F}_{q}^{N}$ are evaluation maps defined by $\widetilde{S}(f)=f(\mathfrak{m})$ and $\widetilde{E}(f)=(f(\mathfrak{m}_{1}),\dots,f(\mathfrak{m}_{N}))$, respectively. We claim that vertical maps in (12) are well-defined and the diagram commutes. This implies that
$\operatorname{Q}_{\mathbb{F}_{q}}(M_{d,\mathbb{F}_{q^{n}}})=\operatorname{Q}_{\mathbb{F}_{q}}(M_{d,\mathcal{O}/\mathfrak{m}})\geq\operatorname{Q}_{\mathbb{F}_{q}}(M_{d,\mathbb{F}_{q}^{N}})=N.$
It is left to prove the claim. We first verify that vertical maps are well-defined. According to Lemma 3.3, we have $v_{\mathfrak{m}}(D)=v_{\mathfrak{m}_{1}}(D)=\cdots=v_{\mathfrak{m}_{N}}(D)=0$. This implies that $v_{\mathfrak{m}}(f)\geq 0$ if $f\in\mathcal{L}(D)$ and $v_{\mathfrak{m}_{j}}(f)\geq 0$ for all $1\leq j\leq N$ if $f\in\mathcal{L}((d-1)D)$, which lead to $\mathcal{L}(D)\subseteq\mathcal{O}$ and $\mathcal{L}((d-1)D)\subseteq\bigcap_{j=1}^{N}\mathcal{O}_{j}$. Therefore, maps $\widetilde{S}\circ E^{-1}$ and $\widetilde{E}\circ S^{-1}$ are well-defined.
Next, we prove the commutativity of (12). For any $a_{1},\dots,a_{d-1}\in\mathbb{F}_{q}^{N}$, we denote $f_{j}\coloneqq E^{-1}(a_{j})\in\mathscr{L}(D)$, $1\leq j\leq d-1$. Thus
@@
$\widetilde{E}\left(\prod_{j=1}^{d-1}f_{j}\right)=\prod_{j=1}^{d-1}\widetilde{E}(f_{j})=\prod_{j=1}^{d-1}a_{i}=\varphi_{d}(a_{1},\dots,a_{d-1}).$
@@
If $d\geq 2$ and $(l-1)\geq(d-1)(n-1)$, then we have $\operatorname{Q}_{\mathbb{F}_{q}}(M_{d,\mathbb{F}_{q^{n}}})\leq\operatorname{Q}_{\mathbb{F}_{q}}(M_{d,\mathbb{F}_{q^{l}}})$.
@@
Since $\mathbb{F}_{q}$ is a perfect field, we observe that (8) (resp. (9)) holds for $(\mathbb{K},\mathbb{F})=(\mathbb{F}_{q},\mathbb{F}_{q^{n}})$ (resp. $(\mathbb{K},\mathbb{F})=(\mathbb{F}_{q},\mathbb{F}_{q^{l}})$). Thus, we have $M_{d,\mathbb{F}_{q^{n}}}\preceq M_{d,n,\mathbb{F}_{q}}\preceq M_{d,\mathbb{F}_{q^{l}}}$, from which we obtain $Q_{\mathbb{F}_{q}}(M_{d,\mathbb{F}_{q^{n}}})\leq Q_{\mathbb{F}_{q}}(M_{d,\mathbb{F}_{q^{l}}})$. Here $M_{d,n,\mathbb{F}_{q}}$ is the tensor corresponding to the multilinear map multiplying (d1)(d-1) $\mathbb{F}_{q}$-polynomials of degree at most (n1)(n-1):
$\Phi_{d}:\underbrace{\mathbb{F}_{q}[x]_{\leq n-1}\times\cdots\times\mathbb{F}_{q}[x]_{\leq n-1}}_{(d-1)\text{~copies}}\to\mathbb{F}_{q}[x]_{(d-1)(n-1)},\quad\Phi_{d}(f_{1},\dots,f_{d-1})=\prod_{j=1}^{d-1}f_{j}.$
@@
For positive integers $m,n$ and $d\geq 2$, we have $\operatorname{Q}_{\mathbb{F}_{q}}(M_{d,\mathbb{F}_{q^{mn}}})\geq\operatorname{Q}_{\mathbb{F}_{q}}(M_{d,\mathbb{F}_{q^{m}}})\operatorname{Q}_{\mathbb{F}_{q^{m}}}(M_{d,\mathbb{F}_{q^{mn}}})$.
@@
$M_{d,\mathbb{F}_{q^{m}}}\in\underbrace{\mathbb{F}_{q^{m}}\otimes_{\mathbb{F}_{q}}\cdots\otimes_{\mathbb{F}_{q}}\mathbb{F}_{q^{m}}}_{d\text{~copies}},\quad M_{d,\mathbb{F}_{q^{mn}}}\in\underbrace{\mathbb{F}_{q^{mn}}\otimes_{\mathbb{F}_{q^{m}}}\cdots\otimes_{\mathbb{F}_{q^{m}}}\mathbb{F}_{q^{mn}}}_{d\text{~copies}}$
@@
$M_{d,\mathbb{F}_{q^{m}}}\boxtimes_{\mathbb{F}^{m}_{q}}M_{d,\mathbb{F}_{q^{mn}}}=M_{d,\mathbb{F}_{q^{mn}}}\in\underbrace{\mathbb{F}_{q^{mn}}\otimes_{\mathbb{F}_{q}}\cdots\otimes_{\mathbb{F}_{q}}\mathbb{F}_{q^{mn}}}_{d\text{~copies}},$
We recall that for tensors $A\preceq B$ and $C\preceq D$, it holds that $A\boxtimes C\preceq B\boxtimes D$. Thus, the desired inequality follows immediately from the definition of the subrank. ∎
@@
For any positive integers $d\geq 2,n$ and prime power $q$, we have $\operatorname{Q}_{\mathbb{F}_{q}}(M_{d,\mathbb{F}_{q^{n}}})\geq\frac{n}{96d^{2}q}$.
@@
We denote $\sigma_{q}(n)\coloneqq\operatorname{Q}_{\mathbb{F}_{q}}(M_{d,\mathbb{F}_{q^{n}}})$ and we first estimate $\sigma_{q^{2}}(n)$. For each positive integer $k$, we let $g_{k}$ and $N_{k}$ (if $k\geq 3$) be positive integers defined as in Lemma A.6. We consider
@@
By (32), we have $N_{m}\geq g_{m}$. The maximality of $m$ and NN leads to